OpenStudy (anonymous):

what is the integral of sqrt(x-1)??

OpenStudy (jamesj):

well, what's this integral: \[ \int x^{1/2} \ dx \] ?

OpenStudy (jamesj):

In general, if \( n \neq -1 \), \[ \int x^n \ dx = \frac{1}{n+1} x^{n+1} + C \]

OpenStudy (jamesj):

So what's the integral I just wrote down?

OpenStudy (jamesj):

Once you know that, use a change of variables, \( u = x - 1 \) and you have the answer to your problem. Following?

OpenStudy (anonymous):

u = x-1 du = dx now you have the integral of u^(1/2) du This gives you 2/3*u^(3/2) Plugging (x-1) back in for u gives you a final solution of: (2/3)*(x-1)^3/2

OpenStudy (anonymous):

+ C after that answer... don't forget that haha